OpenStudy (anonymous):

When finding the line of best fit, a student uses the points (10, 64) and (15, 73), what equation represents their line of best fit?

OpenStudy (anonymous):

A. y = (9/5)x - 46 B. y = (5/9)x + 58 C. y = (5/9)x + 46 D. y = (9/5)x + 46

OpenStudy (anonymous):

I know it's either A or D, I just don't know how to find the slope... but I'm guessing it's D, maybe? I don't think the y-intercept is negative?

OpenStudy (anonymous):

it's D

OpenStudy (anonymous):

Thanks!